What volume of H2 gas at STP would be produced during the passage of 30,000 Coul in the electrolysis of H2 O? 

 



1
Expert's answer
2021-02-28T06:20:49-0500

Ans:-

A coulomb is "1.036 \times10^{-5}" mole of elementary electric charges (electrons, protons, or singly charged ions).

"1" mole of any ideal gas is "22.4 liters" at STP. Hydrogen gas is "H_2" and very close to an ideal gas at STP.


"\Rightarrow" "3.0000\times10^4" "Coulomb" x "1.036 \times10^{-5}" Mole/Coulomb = "3.108 \times10^{-1}" Mole (of H)


"\Rightarrow" "0.3108" Mole (H) "\times" "\dfrac{1\times{H_2}
}{2\times{H}}" "=" "0.1554 mole (H_2)"


"\Rightarrow" "22.4\dfrac{liters}{mole}\times0.1554 mole" "= 3.481 liters" . ("H_2" at STP)